OpenStudy (anonymous):

What is the distance between points (9, 4) and (–3, 4) on a coordinate plane?

OpenStudy (igreen):

Use the distance formula.

OpenStudy (anonymous):

what is that

OpenStudy (igreen):

\(\sf d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (igreen):

(9, 4), (-3, 4) x1 y1 x2 y2

OpenStudy (anonymous):

so I have to complete the fomula

OpenStudy (anonymous):

I am confused

OpenStudy (anonymous):

yes. replace the x1 ,x2 and y1, y2 with the numbers.

OpenStudy (igreen):

Yes

OpenStudy (igreen):

\(\sf d = \sqrt{(-3-9)^2 + (4-4)^2}\)

OpenStudy (anonymous):

yes. now solve it

OpenStudy (igreen):

Oh, well since they have the same y-value, we can just subtract the two numbers. |9 - (-3)|

OpenStudy (anonymous):

how do I solve it

OpenStudy (igreen):

Just subtract |9 - (-3)|

OpenStudy (anonymous):

have you learned subtraction addition and multiplication?

OpenStudy (anonymous):

yes but that thing with all the numbers how do I do that

OpenStudy (anonymous):

igreen has got it down where now all u have to do is subtract 9 from -3

OpenStudy (anonymous):

d= what is that

OpenStudy (anonymous):

dont worry about that part. we are solving for D, as of right now D doesnt matter

OpenStudy (anonymous):

12

OpenStudy (anonymous):

d represents the distance.

OpenStudy (anonymous):

12=9 - -3

OpenStudy (anonymous):

no D= 9-(-3) then when u solve that D=12

OpenStudy (anonymous):

see what i am talking about?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so your distance is 12

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

problema resuelto

OpenStudy (anonymous):

what

OpenStudy (anonymous):

problem solved

OpenStudy (anonymous):

okay